OpenStudy (anonymous):

Find the length of an arc of 40° in a circle with an 8 inch radius. 8/9 inches 16/9 inches 64/9 inches

jigglypuff314 (jigglypuff314):

Arc length (s)= (m/360)(C) where m is the central angle and C is the circumference.C=2 pi*r so = (40/360)( 2 pi (8) ) = (1/9) (16 pi) = 16/9 pi inches

OpenStudy (anonymous):

THANK YOU

jigglypuff314 (jigglypuff314):

you're welcome :)